# Online Convex Optimization with Memory and Limited Predictions

### v1

@@ B.2 Proof of Theorem 2
To proceed, we define the following potential function based on $C_{T}(\cdot)$:
$V_{j}=(1+\gamma)^{j}(C_{T}(\bm{x}^{j})-C_{T}(\bm{x}^{\star})).$ (B.10)
@@
$V_{j+1}-V_{j}$ $=(1+\gamma)^{j}\big{(}(\gamma+1)(C_{T}(\bm{x}^{j+1})-C_{T}(\bm{x}^{j}))+\gamma(C_{T}(\bm{x}^{j})-C_{T}(\bm{x}^{\star}))\big{)}$
$=(1+\gamma)^{j}\big{(}C_{T}(\bm{x}^{j+1})-C_{T}(\bm{x}^{j})+\gamma(C_{T}(\bm{x}^{j+1})-C_{T}(\bm{x}^{\star}))\big{)}.$ (B.11)
To upper bound $V_{j+1}-V_{j}$, we first prove in Appendix D that the following inequality holds for all $j\in\{0,\dots,K-1\}$ and all $\bm{y}\in\mathbb{R}^{dT}$:
@@
Letting $\bm{y}=\bm{x}^{j}$ and $\bm{y}=\bm{x}^{\star}$ in (B.12), we respectively obtain the following:
@@
$C_{T}(\bm{x}^{j+1})-C_{T}(\bm{x}^{\star})$ $\leq\beta^{\prime}\langle\bm{x}^{j}-\bm{x}^{j+1},\bm{x}^{j}-\bm{x}^{\star}\rangle-\frac{\beta^{\prime}}{2}\left\|\bm{x}^{j+1}-\bm{x}^{j}\right\|^{2}-\frac{\mu}{2}\left\|\bm{x}^{j}-\bm{x}^{\star}\right\|^{2}$
$\qquad\qquad\qquad\qquad\qquad\qquad+\langle\nabla C_{T}(\bm{x}^{j})-\bm{g}^{j},\bm{x}^{j+1}-\bm{x}^{\star}\rangle.$ (B.14)
Moreover, for any $\bm{y}=(y_{1},\dots,y_{T})$ with $y_{t}\in\mathcal{X}$ for all $t\in[T]$, we have the following relation proved in Appendix D:
@@
$\mathbb{E}_{\bm{u}^{j}}[V_{j+1}-V_{j}]\leq(1+\gamma)^{j+1}\varepsilon.$ (B.16)
@@
$\mathbb{E}_{\bm{u}^{j}}\big{[}C_{T}(\bm{x}^{j+1})-C_{T}(\bm{x}^{\star})\big{]}\leq\frac{1}{1+\gamma}\big{(}C_{T}(\bm{x}^{j})-C_{T}(\bm{x}^{\star})\big{)}+\varepsilon\quad\forall j\in\{0,\dots,K-1\},$
$\Rightarrow$ $\mathbb{E}_{\mathcal{U}_{K-1}}\big{[}C_{T}(\bm{x}^{K})-C_{T}(\bm{x}^{\star})\big{]}\leq\Big{(}\frac{1}{1+\gamma}\Big{)}^{K}\big{(}C_{T}(\bm{x}^{0})-C_{T}(\bm{x}^{\star})\big{)}+\sum_{j=0}^{K-1}\Big{(}\frac{1}{1+\gamma}\Big{)}^{j}\varepsilon$
$\Rightarrow$ $\mathbb{E}_{\mathcal{U}_{K-1}}\big{[}C_{T}(\bm{x}^{K})-C_{T}(\bm{x}^{\star})\big{]}\leq\Big{(}\frac{1}{1+\gamma}\Big{)}^{K}\big{(}C_{T}(\bm{x}^{0})-C_{T}(\bm{x}^{\star})\big{)}+\frac{\varepsilon}{\gamma},$
where $\mathcal{U}_{K-1}\triangleq(\bm{u}^{0},\dots,\bm{u}^{K-1})$ encapsulates all the randomness of Algorithm 1, completing the proof of the theorem. ∎
@@ Appendix C Omitted Details for the Proof of Theorem 1
C.1 Proof of (B.4)
Recall that $(x_{1}^{\star},\dots,x_{T}^{\star})$ is an optimal solution to (2). Since $\hat{f}_{t}^{c}(\cdot)$ is $\mu$-strongly convex from Lemma 1, we know by the definition that $h_{t}(\cdot)$ is also $\mu$-strongly convex, i.e.,
$h_{t}(x_{t}^{0})-h_{t}(x_{t}^{\star})\leq\nabla h_{t}(x_{t}^{0})^{\top}(x_{t}^{0}-x_{t}^{\star})-\frac{\mu}{2}\left\|x_{t}^{0}-x_{t}^{\star}\right\|^{2}.$ (C.1)
We also have
$\left\|x_{t}^{\star}-x_{t+1}^{0}\right\|^{2}$ $\leq\left\|x_{t}^{\star}-\mathbf{\Pi}_{\mathcal{X}}(x_{t}^{0}-\eta_{t}(\nabla h_{t}(x_{t}^{0})+e_{t}))\right\|^{2}$
$\overset{(a)}{\leq}\left\|x_{t}^{\star}-x_{t}^{0}+\eta_{t}(\nabla h_{t}(x_{t}^{0})+e_{t})\right\|^{2}$
$=\left\|x_{t}^{\star}-x_{t}^{0}\right\|^{2}+2(x_{t}^{\star}-x_{t}^{0})^{\top}\eta_{t}(\nabla h_{t}(x_{t}^{0})+e_{t})+\eta_{t}^{2}\left\|\nabla h_{t}(x_{t}^{0})+e_{t}\right\|^{2},$
where (a) follows from Lemma E.1 in Appendix E. It follows that
$(x_{t}^{0}-x_{t}^{\star})^{\top}\nabla h_{t}(x_{t}^{0})\leq\frac{\left\|x_{t}^{\star}-x_{t}^{0}\right\|^{2}-\left\|x_{t}^{\star}-x_{t+1}^{0}\right\|^{2}}{2\eta_{t}}+(x_{t}^{\star}-x_{t}^{0})^{\top}e_{t}+\frac{\eta_{t}}{2}\left\|\nabla h_{t}(x_{t}^{0})+e_{t}\right\|^{2}.$ (C.2)
Combining (C.1) and (C.2) yields
$h_{t}(x_{t}^{0})-h_{t}(x_{t}^{\star})$ $\leq\frac{\left\|x_{t}^{\star}-x_{t}^{0}\right\|^{2}-\left\|x_{t}^{\star}-x_{t+1}^{0}\right\|^{2}}{2\eta_{t}}+(x_{t}^{\star}-x_{t}^{0})^{\top}e_{t}$
$\qquad+\eta_{t}\left\|\nabla h_{t}(x_{t}^{0})\right\|^{2}+\eta_{t}\left\|e_{t}\right\|^{2}-\frac{\mu}{2}\left\|x_{t}^{0}-x_{t}^{\star}\right\|,$ (C.3)
where we used $\left\|a+b\right\|^{2}\leq 2\left\|a\right\|^{2}+2\left\|b\right\|^{2}$ for $a,b\in\mathbb{R}^{d}$. Meanwhile,
$\left\|x_{t}^{\star}-x_{t+1}^{0}\right\|$ $=\left\|x_{t}^{\star}-x_{t+1}^{\star}+x_{t+1}^{\star}-x_{t+1}^{0}\right\|^{2}$
$=\left\|x_{t}^{\star}-x_{t+1}^{\star}\right\|^{2}+\left\|x_{t+1}^{\star}-x_{t+1}^{0}\right\|^{2}+2(x_{t}^{\star}-x_{t+1}^{\star})^{\top}(x_{t+1}^{\star}-x_{t+1}^{0})$
$\geq\left\|x_{t+1}^{\star}-x_{t+1}^{0}\right\|^{2}-2D\left\|x_{t}^{\star}-x_{t+1}^{\star}\right\|,$ (C.4)
where the inequality follows from Assumption 2 and the fact that $a^{\top}b\geq-\left\|a\right\|\left\|b\right\|$ for $a,b\in\mathbb{R}^{d}$. We then have from (C.3) and (C.4) that
$h_{t}(x_{t}^{0})-h_{t}(x_{t}^{\star})$ $\leq\frac{\left\|x_{t}^{\star}-x_{t}^{0}\right\|^{2}-\left\|x_{t+1}^{\star}-x_{t+1}^{0}\right\|^{2}}{2\eta_{t}}+\frac{D\left\|x_{t}^{\star}-x_{t+1}^{\star}\right\|}{\eta_{t}}$
$\qquad+D\left\|e_{t}\right\|+\eta_{t}\left\|\nabla h_{t}(y_{t})\right\|^{2}+\eta_{t}\left\|e_{t}\right\|^{2}-\frac{\mu}{2}\left\|x_{t}^{0}-x_{t}^{\star}\right\|,$ (C.5)
where we used $(x_{t}^{\star}-x_{t}^{0})^{\top}e_{t}\leq\left\|x_{t}^{\star}-x_{t}^{0}\right\|\left\|e_{t}\right\|\leq D\left\|e_{t}\right\|$. Recalling $\eta_{t}=\frac{1}{t\mu}$, we have
$\sum_{t=1}^{T}\Big{(}\frac{\left\|x_{t}^{\star}-x_{t}^{0}\right\|^{2}-\left\|x_{t+1}^{\star}-x_{t+1}^{0}\right\|^{2}}{2\eta_{t}}-\frac{\mu}{2}\left\|x_{t}^{0}-x_{t}^{\star}\right\|\Big{)}$
$=$ $\sum_{t=1}^{T}(\frac{1}{2\eta_{t}}-\frac{1}{2\eta_{t-1}}-\frac{\mu}{2})\left\|x_{t}^{\star}-x_{t}^{0}\right\|=0,$
[…]
C.2 Proof of (B.5)
@@
$\sum_{t=1}^{T-1}\frac{D\left\|x_{t}^{\star}-x_{t+1}^{\star}\right\|}{\eta_{t}}\overset{(a)}{\leq}\frac{D}{\mu}\sqrt{(\sum_{t=1}^{T-1}\left\|x_{t}^{\star}-x_{t+1}^{\star}\right\|^{2})(\sum_{t=1}^{T-1}\frac{1}{t^{2}})}$
$\overset{(b)}{\leq}$ $\frac{\sqrt{2}D}{\mu}\sqrt{\sum_{t=1}^{T-1}\left\|x_{t}^{\star}-x_{t+1}^{\star}\right\|^{2}}$
$\leq$ $\frac{\sqrt{2}D}{\mu}\sum_{t=1}^{T-1}\left\|x_{t}^{\star}-x_{t+1}^{\star}\right\|,$ (C.6)
@@
To proceed, we upper bound $\left\|e_{t}\right\|$ for all $t\in[T]$, where $e_{t}=e_{t}^{m}+e_{t}^{c}$. We first have from that (B.2)
@@
$\quad|f_{t}(\bar{x}_{t-h+1:t}^{0})-f_{t}^{c}(\bar{x}_{t}^{0})|$
$\leq G\sum_{j=1}^{h-1}\left\|\bar{x}_{t}^{0}-\bar{x}_{t-j}^{0}\right\|$
@@
where (a) uses the relation $x_{t-l+1}^{0}=\Pi_{\mathcal{X}}(x_{t-l}^{0}-\eta_{t-l}(\nabla h_{t-l}(x_{t-l}^{0})+e_{t}))$ and Lemma E.1 in Appendix E.
@@
Combining the upper bounds in (C.6), (C.8), (C.13) and (C.16), noting that $\delta=1/\sqrt{T}\leq 1$ and recalling that $\mathbb{E}[h_{t}(x_{t}^{0})]=\mathbb{E}[\hat{f}_{t}^{c}(x_{t}^{0})]$ and $\mathbb{E}[h_{t}(x_{t}^{\star})]=\mathbb{E}[\hat{f}_{t}^{c}(x_{t}^{\star})]$, we obtain (B.5) from (B.4).
@@ Appendix D Omitted Details in the Proof of Theorem 2
$=-{\beta^{\prime}}^{2}\left\|\bm{x}^{j+1}-\bm{x}^{j}\right\|^{2}+\beta^{\prime}\langle\bm{x}^{j}-\bm{x}^{j+1},\bm{x}^{j}-\bm{y}\rangle+\langle\nabla C_{T}(\bm{x}^{j})-\bm{g}^{j},\bm{x}^{j+1}-\bm{y}\rangle,$
@@
For any $\bm{y}=(y_{1},\dots,y_{T})$ with $y_{t}\in\mathcal{X}$ for all $t\in[T]$, we have
@@
[…]
where the second inequality follows from Assumption 2.
@@
$\overset{(a)}{\leq}D\mathbb{E}_{\bm{u}^{j}}\big{[}\|\nabla C_{T}(\bm{x}^{j})-\widetilde{\bm{g}}^{j}\|\big{]}$
$\leq D\mathbb{E}_{\bm{u}^{j}}\Big{[}\sqrt{\|\nabla C_{T}(\bm{x}^{j})-\widetilde{\bm{g}}^{j}\|^{2}}\Big{]}$
$\overset{(b)}{\leq}D\sqrt{\mathbb{E}_{\bm{u}^{j}}\big{[}\|\nabla C_{T}(\bm{x}^{j})-\widetilde{\bm{g}}^{j}\|^{2}\big{]}}$
$\Rightarrow$ $\frac{1}{D^{2}}\big{(}\mathbb{E}_{\bm{u}^{j}}\big{[}\langle\nabla C_{T}(\bm{x}^{j})-\widetilde{\bm{g}}^{j},\bm{x}^{j+1}-\bm{y}\rangle\big{]}\big{)}^{2}$
@@
where (a) follows from Assumption 2, (b) follows from Jensen’s inequality and the concavity of $\sqrt{\cdot}$ and (c) follows from Lemma 1(b) and Lemma 2(b). [^6] Note the well-known fact that if a differentiable function $f:\mathbb{R}^{d}\to\mathbb{R}$ is convex and $G$-Lipschitz, then $\|\nabla f(x)\|\leq G$ for all $x\in\mathbb{R}^{d}$. Thus, we know from Lemma 1(b) that $\|\nabla C_{T}(\bm{x})\|\leq G\sqrt{Th}$ for all $\bm{x}\in\mathbb{R}^{dT}$. It remains to upper bound the term $-\|C_{T}(\bm{x}^{j})\|^{2}+\mathbb{E}_{\bm{u}^{j}}\big{[}\|\widetilde{\bm{g}}^{j}\|^{2}\big{]}$. Noting that $\sum_{k=s}^{s+h-1}f_{k}(x_{k-h+1:k})$ may be viewed as a function of $x_{s-h+1},\dots,x_{s+h-1}$, denoted as $F_{s}(x_{s-h+1:s+h-1})$, we see from (B.9) that $\widetilde{\bm{g}}^{j}_{s}$ can be equivalently written as
@@
Moreover, using similar arguments to those in the proof of Lemma 1, one can show that $F_{s}(\cdot)$ is $\beta h$-smooth. Now, following similar arguments to those in the proof of [28, Theorem 4], one can show that
@@
[…]
$\leq\frac{D\delta^{\prime}\beta h\sqrt{T}}{2\sqrt{2}(2(2h-1))^{3/4}}+\sqrt{h\beta GT\delta^{\prime}}D(Td+3)^{3/4}+\frac{Dh\sum_{t=1}^{T}\varphi_{t}}{\delta^{\prime}(2(2h-1))^{1/4}}\triangleq\varepsilon.$
@@
$\quad C_{T}(\bm{x}^{j+1})-C_{T}(\bm{x}^{j})+\gamma(C_{T}(\bm{x}^{j+1})-C_{T}(\bm{x}^{\star}))$
$\leq-\frac{\beta^{\prime}}{2}(1+\gamma)\left\|\bm{x}^{j+1}-\bm{x}^{j}\right\|+\gamma\beta^{\prime}\langle\bm{x}^{j}-\bm{x}^{j+1},\bm{x}^{j}-\bm{x}^{\star}\rangle-\frac{\mu\gamma}{2}\left\|\bm{x}^{j}-\bm{x}^{\star}\right\|^{2}+(1+\gamma)\varepsilon$
$\leq-\frac{1}{\beta^{\prime}-\mu}\left\|\beta^{\prime}(\bm{x}^{j+1}-\bm{x}^{j})-\mu(\bm{x}^{j}-\bm{x}^{\star})\right\|^{2}+(1+\gamma)\varepsilon,$

### v3

@@ B.2 Proof of Theorem 2
Let $\bm{x}_{\xi^{\prime}}^{\star}\in\operatorname*{arg\,min}_{\bm{x}\in((1-\xi^{\prime})\mathcal{X})^{T}}C_{T}(\bm{x})$. To proceed, we first establish the convergence to the proxy optimal solution $\bm{x}_{\xi^{\prime}}^{\star}$. Define the potential function:
$W_{j}=(1+\gamma)^{j}(C_{T}(\bm{x}^{j})-C_{T}(\bm{x}_{\xi^{\prime}}^{\star})).$ (B.17)
@@
$W_{j+1}-W_{j}$ $=(1+\gamma)^{j}\big((\gamma+1)(C_{T}(\bm{x}^{j+1})-C_{T}(\bm{x}^{j}))+\gamma(C_{T}(\bm{x}^{j})-C_{T}(\bm{x}_{\xi^{\prime}}^{\star}))\big)$
$=(1+\gamma)^{j}\big(C_{T}(\bm{x}^{j+1})-C_{T}(\bm{x}^{j})+\gamma(C_{T}(\bm{x}^{j+1})-C_{T}(\bm{x}_{\xi^{\prime}}^{\star}))\big).$ (B.18)
To upper bound $W_{j+1}-W_{j}$, we first prove in Appendix D that the following inequality holds for all $j\in\{0,\dots,K-1\}$ and all $\bm{y}\in((1-\xi^{\prime})\mathcal{X})^{T}$:
@@
Letting $\bm{y}=\bm{x}^{j}$ (which holds for $j\geq 1$, and we analyze $j=0$ separately if needed, though the proof structure allows it) and $\bm{y}=\bm{x}_{\xi^{\prime}}^{\star}$ in (B.19) (note that $\bm{x}_{\xi^{\prime}}^{\star}\in((1-\xi^{\prime})\mathcal{X})^{T}$), we respectively obtain the following:
@@
$C_{T}(\bm{x}^{j+1})-C_{T}(\bm{x}_{\xi^{\prime}}^{\star})$ $\leq\beta^{\prime}\langle\bm{x}^{j}-\bm{x}^{j+1},\bm{x}^{j}-\bm{x}_{\xi^{\prime}}^{\star}\rangle-\frac{\beta^{\prime}}{2}\left\|\bm{x}^{j+1}-\bm{x}^{j}\right\|^{2}-\frac{\mu}{2}\left\|\bm{x}^{j}-\bm{x}_{\xi^{\prime}}^{\star}\right\|^{2}$
$\qquad\qquad\qquad\qquad\qquad\qquad+\langle\nabla C_{T}(\bm{x}^{j})-\bm{g}^{j},\bm{x}^{j+1}-\bm{x}_{\xi^{\prime}}^{\star}\rangle.$ (B.21)
Moreover, for any $\bm{y}=(y_{1},\dots,y_{T})$ with $y_{t}\in(1-\xi^{\prime})\mathcal{X}$ for all $t\in[T]$, we have the following relation proved in Appendix D:
@@
$\mathbb{E}_{\bm{u}^{j}}[W_{j+1}-W_{j}]\leq(1+\gamma)^{j+1}\varepsilon.$ (B.23)
@@
$\mathbb{E}_{\bm{u}^{j}}\big[C_{T}(\bm{x}^{j+1})-C_{T}(\bm{x}_{\xi^{\prime}}^{\star})\big]\leq\frac{1}{1+\gamma}\big(C_{T}(\bm{x}^{j})-C_{T}(\bm{x}_{\xi^{\prime}}^{\star})\big)+\varepsilon\quad\forall j\in\{0,\dots,K-1\},$
$\Rightarrow$ $\mathbb{E}_{\mathcal{U}_{K-1}}\big[C_{T}(\bm{x}^{K})-C_{T}(\bm{x}_{\xi^{\prime}}^{\star})\big]\leq\Big(\frac{1}{1+\gamma}\Big)^{K}\big(C_{T}(\bm{x}^{0})-C_{T}(\bm{x}_{\xi^{\prime}}^{\star})\big)+\sum_{j=0}^{K-1}\Big(\frac{1}{1+\gamma}\Big)^{j}\varepsilon$
$\Rightarrow$ $\mathbb{E}_{\mathcal{U}_{K-1}}\big[C_{T}(\bm{x}^{K})-C_{T}(\bm{x}_{\xi^{\prime}}^{\star})\big]\leq\Big(\frac{1}{1+\gamma}\Big)^{K}\mathbb{E}\big[C_{T}(\bm{x}^{0})-C_{T}(\bm{x}_{\xi^{\prime}}^{\star})\big]+\frac{\varepsilon}{\gamma},$
where $\mathcal{U}_{K-1}\triangleq(\bm{u}^{0},\dots,\bm{u}^{K-1})$ encapsulates all the randomness of Algorithm 1.
Overall Bound on the Regret $R(\mathcal{A})$. We now combine this result with the initialization regret to bound the total dynamic regret against the true optimal $\bm{x}^{\star}$. We decompose the regret as:
$\mathbb{E}[R(\mathcal{A})]=\mathbb{E}\big[C_{T}(\bm{x}^{K})-C_{T}(\bm{x}^{\star})\big]=\underbrace{\mathbb{E}\big[C_{T}(\bm{x}^{K})-C_{T}(\bm{x}_{\xi^{\prime}}^{\star})\big]}_{\text{Term A}}+\underbrace{\Big(C_{T}(\bm{x}_{\xi^{\prime}}^{\star})-C_{T}(\bm{x}^{\star})\Big)}_{\text{Term B}}.$
Bounding Term B (Approximation Error): Since $\bm{x}^{\star}$ is in $\mathcal{X}^{T}$ and contains the origin, $(1-\xi^{\prime})\bm{x}^{\star}\in((1-\xi^{\prime})\mathcal{X})^{T}$. The approximation error is bounded by:
Term B $=C_{T}(\bm{x}_{\xi^{\prime}}^{\star})-C_{T}(\bm{x}^{\star})$
$\overset{\text{(a)}}{\leq}C_{T}((1-\xi^{\prime})\bm{x}^{\star})-C_{T}(\bm{x}^{\star})$
$\overset{\text{(b)}}{\leq}\xi^{\prime}GT\sqrt{h}D,$
where (a) is due to $\bm{x}_{\xi^{\prime}}^{\star}\in\operatorname*{arg\,min}_{\bm{x}\in((1-\xi^{\prime})\mathcal{X})^{T}}C_{T}(\bm{x})$; (b) follows the same proof logic as (B.15).
Bounding Term A and Combining: We substitute the bound for Term A derived before:
$\text{Term A}\leq\Big(\frac{1}{1+\gamma}\Big)^{K}\mathbb{E}\big[C_{T}(\bm{x}^{0})-C_{T}(\bm{x}_{\xi^{\prime}}^{\star})\big]+\frac{\varepsilon}{\gamma}.$
We bound the initial error term $\mathbb{E}\big[C_{T}(\bm{x}^{0})-C_{T}(\bm{x}_{\xi^{\prime}}^{\star})\big]$ by relating it to the total initialization regret. Since $C_{T}(\bm{x}_{\xi^{\prime}}^{\star})\geq C_{T}(\bm{x}^{\star})$, we have:
$\mathbb{E}\big[C_{T}(\bm{x}^{0})-C_{T}(\bm{x}_{\xi^{\prime}}^{\star})\big]$ $=\mathbb{E}\big[C_{T}(\bm{x}^{0})-C_{T}(\bm{x}^{\star})\big]-\big(C_{T}(\bm{x}_{\xi^{\prime}}^{\star})-C_{T}(\bm{x}^{\star})\big)$
$\leq\mathbb{E}[R(\mathcal{A}_{\text{init}})].$
Substituting these back into the regret decomposition, we obtain the final result:
$\mathbb{E}[R(\mathcal{A})]\leq\Big(\frac{1}{1+\gamma}\Big)^{K}\mathbb{E}[R(\mathcal{A}_{\text{init}})]+\frac{\varepsilon}{\gamma}+\xi^{\prime}GT\sqrt{h}D.$
This completes the proof of Theorem 2.
@@ Appendix C Omitted Details for the Proof of Theorem 1
C.1 Proof of the bounds of the four terms in (B.12)
@@
$\sum_{t=1}^{T-1}\frac{(1-\xi)D\left\|\hat{x}_{t}^{\star}-\hat{x}_{t+1}^{\star}\right\|}{\eta_{t}}\overset{(a)}{\leq}\frac{(1-\xi)D}{\mu}\sqrt{(\sum_{t=1}^{T-1}\left\|\hat{x}_{t}^{\star}-\hat{x}_{t+1}^{\star}\right\|^{2})(\sum_{t=1}^{T-1}\frac{1}{t^{2}})}$
$\overset{(b)}{\leq}$ $\frac{\sqrt{2}(1-\xi)D}{\mu}\sqrt{\sum_{t=1}^{T-1}\left\|\hat{x}_{t}^{\star}-\hat{x}_{t+1}^{\star}\right\|^{2}}$
$\leq$ $\frac{\sqrt{2}(1-\xi)D}{\mu}\sum_{t=1}^{T-1}\left\|\hat{x}_{t}^{\star}-\hat{x}_{t+1}^{\star}\right\|=\frac{\sqrt{2}(1-\xi)^{2}D}{\mu}V_{T},$ (C.1)
@@
To proceed, we upper bound $\left\|e_{t}\right\|$ for all $t\in[T]$, where $e_{t}=e_{t}^{m}+e_{t}^{c}$. We first have from (B.3) that
@@
$\quad|f_{t}(\bar{x}_{t-h+1:t}^{0})-f_{t}^{c}(\bar{x}_{t}^{0})|$ $\leq G\sum_{j=1}^{h-1}\left\|\bar{x}_{t}^{0}-\bar{x}_{t-j}^{0}\right\|$
@@
where (a) uses the relation $x_{t-l+1}^{0}=\Pi_{(1-\xi)\mathcal{X}}(x_{t-l}^{0}-\eta_{t-l}(\nabla h_{t-l}(x_{t-l}^{0})+e_{t}))$ and Lemma E.1 in Appendix E.
@@
[…]
@@ Appendix D Omitted Details in the Proof of Theorem 2
$=-\beta^{\prime}\left\|\bm{x}^{j+1}-\bm{x}^{j}\right\|^{2}+\beta^{\prime}\langle\bm{x}^{j}-\bm{x}^{j+1},\bm{x}^{j}-\bm{y}\rangle+\langle\nabla C_{T}(\bm{x}^{j})-\bm{g}^{j},\bm{x}^{j+1}-\bm{y}\rangle,$
@@
For any $\bm{y}=(y_{1},\dots,y_{T})$ with $y_{t}\in(1-\xi^{\prime})\mathcal{X}$ for all $t\in[T]$, we have
@@
$\text{(ii)}\leq\mathbb{E}_{\bm{u}^{j}}\big[\|\bm{g}^{j}-\widetilde{\bm{g}}^{j}\|\|\bm{x}^{j+1}-\bm{y}\|\big]\leq\frac{(1-\xi^{\prime})Dh\sum_{t=1}^{T}\varphi_{t}}{\delta^{\prime}(2(2h-1))^{1/4}},$
where the second inequality follows from Remark 1 (noting $\bm{x}^{j+1},\bm{y}\in((1-\xi^{\prime})\mathcal{X})^{T}$).
@@
$\overset{(a)}{\leq}(1-\xi^{\prime})D\mathbb{E}_{\bm{u}^{j}}\big[\|\nabla C_{T}(\bm{x}^{j})-\widetilde{\bm{g}}^{j}\|\big]$
$\leq(1-\xi^{\prime})D\mathbb{E}_{\bm{u}^{j}}\Big[\sqrt{\|\nabla C_{T}(\bm{x}^{j})-\widetilde{\bm{g}}^{j}\|^{2}}\Big]$
$\overset{(b)}{\leq}(1-\xi^{\prime})D\sqrt{\mathbb{E}_{\bm{u}^{j}}\big[\|\nabla C_{T}(\bm{x}^{j})-\widetilde{\bm{g}}^{j}\|^{2}\big]}$
$\Rightarrow$ $\frac{1}{(1-\xi^{\prime})^{2}D^{2}}\big(\mathbb{E}_{\bm{u}^{j}}\big[\langle\nabla C_{T}(\bm{x}^{j})-\widetilde{\bm{g}}^{j},\bm{x}^{j+1}-\bm{y}\rangle\big]\big)^{2}$
@@
where (a) follows from Remark 1, (b) follows from Jensen’s inequality and the concavity of $\sqrt{\cdot}$ and (c) follows from Lemma 1(b) and Lemma 2(b). [^6] Note the well-known fact that if a differentiable function $f:\mathbb{R}^{d}\to\mathbb{R}$ is convex and $G$-Lipschitz, then $\|\nabla f(x)\|\leq G$ for all $x\in\mathbb{R}^{d}$. Thus, we know from Lemma 1(b) that $\|\nabla C_{T}(\bm{x})\|\leq G\sqrt{Th}$ for all $\bm{x}\in\mathbb{R}^{dT}$. It remains to upper bound the term $\|C_{T}(\bm{x}^{j})\|^{2}+\mathbb{E}_{\bm{u}^{j}}\big[\|\widetilde{\bm{g}}^{j}\|^{2}\big]$. Noting that $\sum_{k=s}^{s+h-1}f_{k}(x_{k-h+1:k})$ may be viewed as a function of $x_{s-h+1},\dots,x_{s+h-1}$, denoted as $F_{s}(x_{s-h+1:s+h-1})$, we see from (B.16) that $\widetilde{\bm{g}}^{j}_{s}$ can be equivalently written as
@@
Moreover, using similar arguments to those in the proof of Lemma 1, one can show that $F_{s}(\cdot)$ is $\beta h$-smooth. Now, following similar arguments to those in the proof of [nesterov2017random, Theorem 4], one can show that
@@
Thus, we can upper bound $\text{(i)}+\text{(ii)}$ as
$\text{(i)}+$ $\text{(ii)}\leq\sqrt{\frac{(1-\xi^{\prime})^{2}D^{2}\delta^{\prime 2}\beta^{2}h^{2}T}{8(2(2h-1))^{3/2}}+(1-\xi^{\prime})^{2}D^{2}Gh\beta T(Td+3)^{3/2}\delta^{\prime}}+\frac{(1-\xi^{\prime})Dh\sum_{t=1}^{T}\varphi_{t}}{\delta^{\prime}(2(2h-1))^{1/4}}$
$\leq\frac{(1-\xi^{\prime})D\delta^{\prime}\beta h\sqrt{T}}{2\sqrt{2}(2(2h-1))^{3/4}}+\sqrt{h\beta GT\delta^{\prime}}(1-\xi^{\prime})D(Td+3)^{3/4}+\frac{(1-\xi^{\prime})Dh\sum_{t=1}^{T}\varphi_{t}}{\delta^{\prime}(2(2h-1))^{1/4}}\triangleq\varepsilon.$
@@
$\quad C_{T}(\bm{x}^{j+1})-C_{T}(\bm{x}^{j})+\gamma(C_{T}(\bm{x}^{j+1})-C_{T}(\bm{x}_{\xi^{\prime}}^{\star}))$
$\leq-\frac{\beta^{\prime}}{2}(1+\gamma)\left\|\bm{x}^{j+1}-\bm{x}^{j}\right\|+\gamma\beta^{\prime}\langle\bm{x}^{j}-\bm{x}^{j+1},\bm{x}^{j}-\bm{x}_{\xi^{\prime}}^{\star}\rangle-\frac{\mu\gamma}{2}\left\|\bm{x}^{j}-\bm{x}_{\xi^{\prime}}^{\star}\right\|^{2}+(1+\gamma)\varepsilon$
$\leq-\frac{1}{\beta^{\prime}-\mu}\left\|\beta^{\prime}(\bm{x}^{j+1}-\bm{x}^{j})-\mu(\bm{x}^{j}-\bm{x}_{\xi^{\prime}}^{\star})\right\|^{2}+(1+\gamma)\varepsilon\leq(1+\gamma)\varepsilon,$